OpenStudy (anonymous):

i need help factoring c^2 - 6cd + 9d^2 -4 Can you please write it out so i can see how it is done

jhonyy9 (jhonyy9):

c^2 - 2*c*3d +(3d)^2 so after this help i think that will be easy

OpenStudy (anonymous):

as difference of squares

OpenStudy (anonymous):

can u draw it out so i can understand it beter

jhonyy9 (jhonyy9):

yes but firstly you need to make what i have wrote after this we can using (-4) to ending this exercise

OpenStudy (anonymous):

i dnt understand what u wrote out

jhonyy9 (jhonyy9):

so for you can understanding better look this please a*a - 2*a*b +b*b = ???

jhonyy9 (jhonyy9):

so firstly i like to make you learning this but if you wann i can solving it shortly

OpenStudy (anonymous):

Can u draw it out so i can get a visual

jhonyy9 (jhonyy9):

how many is (a-b)^2 = ???

jhonyy9 (jhonyy9):

so equal a^2 -2ab +b^2 = a*a -2ab +b*b yes ?

OpenStudy (anonymous):

yes i get that

jhonyy9 (jhonyy9):

can you make it coresponding with what i have wrote there firstly ?

OpenStudy (anonymous):

no i dont get how to solve it thats why i am asking for someone to draw it on the board so i can get whats going on

jhonyy9 (jhonyy9):

(c-3d)^2 -4 =

jhonyy9 (jhonyy9):

this is clearly ?

jhonyy9 (jhonyy9):

and now you can using the formula a^2 -b^2 =(a-b)(a+b)

OpenStudy (anonymous):

thats just it i dont know how to put it in the (a-b)(a+b) form

jhonyy9 (jhonyy9):

so from c^2 -6cd +9d^2 we have got (c-3d)^2 what will be in this last point a and 4=2*2 so hence 2^2 will be b^2 and now you just need to put these inside formula

jhonyy9 (jhonyy9):

ok ?

jhonyy9 (jhonyy9):

(c-3d)^2 will be a^2

jhonyy9 (jhonyy9):

than there we will get (c-3d)^2 -4=(c-3d-2)(c-3d+2)